OpenStudy (peterg2001):

help please need help fast 1# To solve -9 - 16m = 7, what steps would you use? 9 -16 2# To solve -3 = -19 + d/5 what steps would you use? 19 5 3#Which of the following equations could you solve by first adding six and then dividing by negative three? 3x - 6 = -9 6 - 3x = -9 -3x - 6 = -9 x/-3 + 6 = -9 4# Which of the following is a solution of the equation 9 - 2x = -1? -5 5 4 -4

OpenStudy (peterg2001):

@silverr @ThisIsAdriAtHerBest

OpenStudy (anonymous):

#.1 -9 - 16m = 7 7 + -9 = ? ? / 16 = m

OpenStudy (anonymous):

#2. -3 = -19 + d/5 -3 - (-19) = ? ? / 5 = d

OpenStudy (anonymous):

3 # Which of the following equations could you solve by first adding six and then dividing by negative three? 3x - 6 = -9

OpenStudy (anonymous):

#4. Which of the following is a solution of the equation 9 - 2x = -1? -1 + 9 = ? ? / 2 = x

OpenStudy (anonymous):

@peterg2001 , can you solve to get the answers?